The angle between two vectors, u and v, is 30 degrees and the dot product is 24. If the magnitude of vector u is √12, then calculate the magnitude of vector v.
Answer by amarjeeth123(569)

The dot product is defined as u.v=uvcos(theta)
Here theta=30 degrees
u=√12
Dot product is 24.
cos30=√3/2
Plugging in the values we get,
24=√12(√3/2)v
24=√36/2(v)
24=3v
v=8
The magnitude of vector v is 8 units.
Answer=8 units.